OpenStudy (anonymous):

Find an equation of the line containing the given pair of points. (1/4, -1/2)and (3/4, 6) Type your answer in slope intercept form.

OpenStudy (anonymous):

First find the slope(m) by using y2-y1/x2-x1 then for the equation use this formula y-y1=m(x-x1)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

6.5/0.5 is the slope but does this reduce

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ \frac{1}{4}}}\quad ,&{\color{blue}{ -\frac{1}{2}}})\quad &({\color{red}{ \frac{3}{4}}}\quad ,&{\color{blue}{ 6}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

1/4=x1 and 3/4=x2 -1/2=y1 and 6=y2

OpenStudy (anonymous):

ok

OpenStudy (jdoe0001):

hmm lemme fix that a bit

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ \frac{1}{4}}}\quad ,&{\color{blue}{ -\frac{1}{2}}})\quad &({\color{red}{ \frac{3}{4}}}\quad ,&{\color{blue}{ 6}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{6}}-{\color{blue}{ \left(-\frac{1}{2}\right)}}}{{\color{red}{ \frac{3}{4}}}-{\color{red}{ \frac{1}{4}}}}\implies \cfrac{6+\frac{1}{2}}{\frac{3}{4}-\frac{1}{4}} \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form} \)

OpenStudy (jdoe0001):

har har saw that :P is actually red :) but orange sounds tastier